A manufacturer knows that if x goods are demanded on a particular week, the total cost and revenue functions will be:𝐢(π‘₯) = 14 + 3π‘₯ π‘Žπ‘›π‘‘ 𝑅(π‘₯) = 18π‘₯ βˆ’ 2π‘₯ 2 respectively.

i. Calculate the level of demand that will maximize profits (6 marks)

ii. Calculate the amount of profit that will be realized at this maximum point. (4 marks)Β 


1
Expert's answer
2021-11-24T02:33:39-0500

i)

profit function:

"P(x)=\ud835\udc45(\ud835\udc65)-\ud835\udc36(\ud835\udc65)=18x-2x^2-14-3x=-2x^2+15x-14"

"P'(x)=-4x+15=0"

level of demand that will maximize profits:

"x=15\/4=3.75"


ii)

amount of profit that will be realized at this maximum point:

"P_{max}(x)=-2\cdot3.75^2+15\cdot3.75-14=14.125"